### v1

# Conditional Dependence via U-Statistics Pruning

Abstract
The problem of measuring conditional dependence between two random phenomena arises when a third one (a confounder) has a potential influence on the amount of information shared by the original pair. A typical issue in this challenging problem is the inversion of ill-conditioned autocorrelation matrices. This paper presents a novel measure of conditional dependence based on the use of incomplete unbiased statistics of degree two, which allows to re-interpret independence as uncorrelatedness on a finite-dimensional feature space. This formulation enables to prune data according to the observations of the confounder itself, thus avoiding matrix inversions altogether. Moreover, the proposed approach is articulated as an extension of the Hilbert-Schmidt independence criterion, which becomes expressible through kernels that operate on 4-tuples of data.
Index Terms:
Hilbert-Schmidt independence criterion (HSIC), kernel methods, co-information, U-Statistics, conditional dependence.
I Introduction
The problem of measuring statistical dependence emerges in many fields, such as multivariate statistical analysis [1, Ch. 6], wireless sensor networks [2, Ch. 2], data analytics [3], and statistical array processing [4]. The Hilbert-Schmidt independence criterion (HSIC[5] provides a versatile non-negative measure of statistical dependence that becomes null if and only if the data pair is independent. Grounded in the theory of reproducing kernel Hilbert space s (RKHS), the HSIC succeeds in discovering both linear and nonlinear relationships in data. A well-known interpretation of the HSIC is that dependence produces a correlation of distances [6], i.e. close couples in one data set coincide with close couples in the other data set. Another insightful interpretation is that the HSIC measures statistical dependence by gauging the correlation of data implicitly mapped onto an infinite-dimensional space. This basic idea of translating dependence to correlation also admits a more traditional implementation by explicitly mapping data onto a finite-dimensional space [7], with the advantage of leveraging classical second-order statistics for inferring information-theoretic measures.
Conditional dependence becomes relevant when a third phenomenon might explain, mediate, or confound the apparent relationship exhibited by the original random pair, which happens in fields such as causal discovery and Bayesian network learning [8]. Although mapping data onto a space of higher dimension is still useful for that purpose, it is still necessary to measure conditional correlation in this newfound space. This method usually leads to the computation of the Schur complement of high-dimensional matrices [9, Ch. 4], which is known to have numerical issues concerned with matrix inversions and regularization due to the usually low rank structure of the involved autocorrelation matrices [10].
This letter presents a procedure for measuring conditional dependence by statistically conditioning the observed data to a potential confounder, which bypasses the matrix inversion problems. First, a novel theoretical derivation of the classical HSIC is provided in Sec. II, by translating the problem of statistical dependence into one of correlation after mapping the data onto finite-dimensional spaces based on steering vectors [7]. Sec. III briefly reviews the theory of unbiased statistics (U-Statistics) [11], which is then employed in Sec. IV to show that conditional dependence can be accomplished by just pruning U-Statistics under the control of the confounder [12]. The obtained measure, the conditional HSIC (C-HSIC), is based on a signal processing structure resembling kernel methods, which embraces the HSIC as a particular case when the U-Statistic is complete. However, it has the distinctive feature of being articulated on 4-tuples instead of the classical processing based on pairs of data from each phenomenon.
II Marginal dependence as correlation
Let $\mathbf{d}:\mathbb{R}\to\mathbb{C}^{M}$ be a mapping based on windowed steering vectors. The $m$th element of $\mathbf{d}$ can be expressed as
[$\mathbf{d}(\bm{\cdot})]_{m}\triangleq\tfrac{1}{\sqrt[4]{M}}\mathrm{G}\bigl{(}\tfrac{m}{\sqrt{M}}\bigr{)}\exp\bigl{(}\mathrm{j}\pi\tfrac{m}{\sqrt{M}}\bm{\cdot}\bigr{)},$ (1)
where $m\in\{-M/2,\dots,M/2-1\}$ and $\mathrm{G}:\mathbb{R}\to\mathbb{R}$ is an even, absolutely integrable function with unit $L^{2}$-norm and maximum at $\mathrm{G}(0)$. Given two random variables $\mathsf{x}$ and $\mathsf{y}$, we define a pair of transformed random vectors [^1] Without loss of generality, and for the sake of simplicity, we let $\bm{\mathsf{u}}$ and $\bm{\mathsf{v}}$ have the same dimensionality. $\bm{\mathsf{u}}\triangleq\mathbf{d}(\mathsf{x})$ and $\bm{\mathsf{v}}\triangleq\mathbf{d}(\mathsf{y})$. Their cross-covariance matrix $\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\in\mathbb{C}^{M\times M}$ is defined as $\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\triangleq\operatorname{\mathrm{E}}\bigl{[}\bm{\mathsf{u}}\bm{\mathsf{v}}^{\mathrm{H}}\bigr{]}-\operatorname{\mathrm{E}}[\bm{\mathsf{u}}]\operatorname{\mathrm{E}}[\bm{\mathsf{v}}]^{\mathrm{H}}$, where $\operatorname{\mathrm{E}}[\bm{\cdot}]$ denotes the expectation operator. Then, the following implication holds [5]:
$\lim_{M\to\infty}\lVert\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\rVert_{\mathrm{F}}^{2}=0\quad\iff\quad\mathsf{x}\;{\perp\!\!\!\perp}\;\mathsf{y}\;,$ (2)
[…]
Consider $L$ i.i.d. samples of $\mathsf{x}$ and $\mathsf{y}$, $\{x(l),y(l)\}_{l=1,\dotsc,L}$. From them, we obtain the two transformed complex data matrices $\mathbf{U}\in\mathbb{C}^{M\times L}$ and $\mathbf{V}\in\mathbb{C}^{M\times L}$, defined as follows:
$\mathbf{U}$ $\triangleq[\mathbf{u}_{1},\dotsc,\mathbf{u}_{L}]=[\mathbf{d}(x(1)),\dotsc,\mathbf{d}(x(L))]$ (3a)
$\mathbf{V}$ $\triangleq[\mathbf{v}_{1},\dotsc,\mathbf{v}_{L}]=[\mathbf{d}(y(1)),\dotsc,\mathbf{d}(y(L))].$ (3b)
The unbiased sample cross-covariance between the transformed vectors is given by
$\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ $\triangleq\tfrac{1}{L-1}\sum_{l=1}^{L}\Biggl{(}\mathbf{u}_{l}-\tfrac{1}{L}\sum_{i=1}^{L}\mathbf{u}_{i}\Biggr{)}\Biggl{(}\mathbf{v}_{l}-\tfrac{1}{L}\sum_{j=1}^{L}\mathbf{v}_{j}\Biggr{)}^{\mathrm{H}}$
$=\tfrac{1}{L-1}\mathbf{U}\mathbf{P}\mathbf{V}^{\mathrm{H}},\quad$ (4)
where $\mathbf{P}\triangleq\mathbf{I}-\tfrac{1}{L}\mathbf{1}\mathbf{1}^{\mathrm{T}}\in\mathbb{R}^{L\times L}$. Using (2), a marginal dependence measure is given by
$\lVert\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}\rVert_{\mathrm{F}}^{2}=\operatorname{\mathrm{tr}}(\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}^{\mathrm{H}}\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}})$ $=\operatorname{\mathrm{tr}}\mathopen{}\mathclose{{}\left(\tfrac{1}{(L-1)^{2}}\mathbf{V}\mathbf{P}^{\mathrm{T}}\mathbf{U}^{\mathrm{H}}\mathbf{U}\mathbf{P}\mathbf{V}^{\mathrm{H}}}\right)$
$=\tfrac{1}{(L-1)^{2}}\operatorname{\mathrm{tr}}(\mathbf{P}\mathbf{U}^{\mathrm{H}}\mathbf{U}\mathbf{P}\mathbf{V}^{\mathrm{H}}\mathbf{V}),$ (5)
where the circularity of the trace operator has been used. To see the link with the HSIC, let us examine the limit for $M\to\infty$ of the Gramm (kernel) matrices $\mathbf{K}=\lim_{M\rightarrow\infty}\mathbf{U}^{\mathrm{H}}\mathbf{U}\in\mathbb{R}^{L\times L}$ and $\mathbf{Q}=\lim_{M\rightarrow\infty}\mathbf{V}^{\mathrm{H}}\mathbf{V}\in\mathbb{R}^{L\times L}$. The elements of $\mathbf{K}$ (and analogously those of $\mathbf{Q}$) are the following:
$[\mathbf{K}]_{l,l^{\prime}}$ $=\lim_{M\to\infty}\mathbf{d}^{\mathrm{H}}(x(l))\mathbf{d}(x(l^{\prime}))=$
$\lim_{M\to\infty}$ $\sum_{m=-\frac{M}{2}}^{\frac{M}{2}-1}\tfrac{1}{\sqrt{M}}\mathrm{G}^{2}\bigl{(}\tfrac{m}{\sqrt{M}}\bigr{)}\exp\biggl{(}\mathrm{j}\frac{2\pi(x(l^{\prime})-x(l))m}{\sqrt{M}}\biggr{)}$ (6a)
$=\int_{-\infty}^{\infty}\mathrm{G}^{2}(f)\exp(\mathrm{j}2\pi(x(l^{\prime})-x(l))f)\operatorname{\mathrm{d}\!}f,$ (6b)
where (6b) is the result of looking at (6a) as a Darboux sum. Then, given (II) and taking the limit $M\to\infty$, we finally obtain the HSIC [5, Sec. 3.1]:
[…]
with $\mathbf{x}=[x(1),\dotsc,x(L)]^{\mathrm{T}}$ and $\mathbf{y}=[y(1),\dotsc,y(L)]^{\mathrm{T}}$.
As an alternative to RKHS theory, the result in (7) has been derived here by means of the finite-dimensional universal mapping proposed in (1), which will play a fundamental role in the extension to the conditional case proposed later on. By defining $\kappa(\bm{\cdot})$ as the inverse Fourier transform222 From the properties imposed on $\mathrm{G}^{2}(f)$, $\kappa(\bm{\cdot})$ becomes naturally an autocorrelation translation-invariant kernel, such that $\kappa(0)=1\geq|\kappa(s)|$ and $\kappa(\infty)=\kappa(-\infty)=0$. The reader is referred to [15, Sec. 1.4] where these ideas emerge in the light of Bochner’s theorem. of $\mathrm{G}^{2}(f)$, the $(l,l^{\prime})$ entries of $\mathbf{K}$ and $\mathbf{Q}$ are just the evaluation of $\kappa(\bm{\cdot})$ at a value given by the difference between two data samples:
$[\mathbf{K}]_{l,l^{\prime}}=\kappa\bigl{(}x(l^{\prime})-x(l)\bigr{)}.$ (8)
Motivated by this structure, let us introduce a more compact and general notation that will become relevant later on
$[\mathbf{S}(\mathbf{a},\mathbf{b})]_{l,l^{\prime}}\triangleq\kappa\bigl{(}[\mathbf{b}]_{l}-[\mathbf{a}]_{l^{\prime}}\bigr{)},$ (9)
being $\mathbf{S}(\mathbf{a},\mathbf{b})$ a general kernel matrix obtained from two generic vectors $\mathbf{a}$ and $\mathbf{b}$. With it, we can compactly write both kernel matrices using (9) as follows:
$\mathbf{K}=\mathbf{S}(\mathbf{x},\mathbf{x}),\quad\mathbf{Q}=\mathbf{S}\mathopen{}\mathclose{{}\left(\mathbf{y},\mathbf{y}}\right).$ (10)
In summary, the HSIC results from measuring correlation in a finite-dimensional space based on steering vectors. From that, kernel formulations arise in a second stage once dimensionality grows without bound.
III Sample covariance matrix as an incomplete U-Statistic
Next we present the U-Statistic reformulation jointly with its incomplete expression. The objective is to provide an alternative covariance matrix estimation based on pairwise differences between the original samples that will serve as an introduction to the pruning of data pairs.
Consider a list containing all the unique $K_{\max}\triangleq L(L-1)/2$ pairs that can be constructed from $L$ different samples of $\mathsf{x}$ or $\mathsf{y}$. This list is ordered arbitrarily such that a single index $k$ identifies both elements of a pair:
$k\mapsto(x(\mathrm{f}_{1}(k)),x(\mathrm{f}_{2}(k))),\quad k\mapsto(y(\mathrm{f}_{1}(k)),y(\mathrm{f}_{2}(k))).$ (11)
Functions $\mathrm{f}_{1}(\bm{\cdot})$ and $\mathrm{f}_{2}(\bm{\cdot})$ return the corresponding indices of each term of the pair. Given $K\leq K_{\max}$ indices, we construct the pairwise differences
$\mathring{\mathbf{u}}_{k}\triangleq\frac{1}{\sqrt{2}}\mathopen{}\mathclose{{}\left(\mathbf{u}_{f_{1}(k)}-\mathbf{u}_{f_{2}(k)}}\right),\quad\mathring{\mathbf{v}}_{k}\triangleq\frac{1}{\sqrt{2}}\mathopen{}\mathclose{{}\left(\mathbf{v}_{f_{1}(k)}-\mathbf{v}_{f_{2}(k)}}\right),$ (12)
for $k\in\{1,\dots,K\}$, corresponding to the samples of the new zero-mean virtual sources $\mathring{\bm{\mathsf{u}}}$ and $\mathring{\bm{\mathsf{v}}}$. Their sample cross-covariance matrix $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}\in\mathbb{C}^{M\times M}$ is then the following:
$\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}=\tfrac{1}{K}\mathring{\mathbf{U}}\mathring{\mathbf{V}}^{\mathrm{H}},$ (13)
where $\mathring{\mathbf{U}}\triangleq[\mathring{\mathbf{u}}_{1},\dots,\mathring{\mathbf{u}}_{K}]\in\mathbb{C}^{M\times K}$ and $\mathring{\mathbf{V}}\triangleq[\mathring{\mathbf{v}}_{1},\dots,\mathring{\mathbf{v}}_{K}]\in\mathbb{C}^{M\times K}$. Note that, thanks to the constant factor in (12), $\mathring{\bm{\mathsf{u}}}$ and $\mathring{\bm{\mathsf{v}}}$ have the same average covariance matrix as $\bm{\mathsf{u}}$ and $\bm{\mathsf{v}}$, respectively. In contrast to (4), $\mathbf{P}$ is missing in (13) as a result of constructing zero-mean virtual data in (12), which will lead to a cleaner implementation and fewer matrix operations.
It is worth noting that expression (13) is in fact an instance of a U-Statistic. For $K=K_{\text{max}}$, i.e. when all data pairs are taken, we obtain the equality $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}=\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ from U-Statistics theory [11]. In contrast, for $K<K_{\text{max}}$, i.e. when (13) is an incomplete U-Statistic, although $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}$ remains unbiased, its estimation variance increases due to the pruning of data. Nevertheless, $\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}}$ is still a consistent estimate of $\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ provided that $K\rightarrow\infty$ as $L\rightarrow\infty$.
Remark 1 (Robustness to pruning)
A notable property of incomplete U-statistics is that their robustness against data pruning increases the larger $L$ is [12]. To briefly illustrate this property, consider taking only the $K=\mathopen{}\mathclose{{}\left\lfloor L/2}\right\rfloor$ data pairs with no indices in common. The number of remaining, unused, data pairs is equal to $L(L-1)/2-\lfloor L/2\rfloor$, which increases with $O(L^{2})$. To only use the i.i.d., nonrepeated, terms is effectively equivalent to computing $\widehat{\mathbf{C}}_{\bm{\mathsf{u}},\bm{\mathsf{v}}}$ with half of the available samples [16]. It is then safe to assume that their contribution to the overall accuracy of the sample covariance is higher than those with repeated indices. Therefore, the larger $L$ is, the higher the amount of pairs that can be pruned for some specified degradation in the estimation accuracy of the resulting sample covariance. The implication is that $K$ in the incomplete U-Statistic can be designed to grow with $O(L)$ instead of $O(L^{2})$, which provides a lot of flexibility for pruning, and will be used for choosing the number of pairs in the next section.
IV Conditional dependence via U-Statistics
The conditional cross-covariance matrix between $\bm{\mathsf{u}}$ and $\bm{\mathsf{v}}$ is defined as:
$\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}}\triangleq\int_{\mathbb{R}}\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}=z}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z)=\int_{\mathbb{R}}\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}=z}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z)$ (14)
where $\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}=z}=\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}=z}$ (given (12)) is the cross-covariance matrix conditioned to a specific value $z$ of a confounder $\mathsf{z}$, and $F_{\mathsf{z}}(z)$ is its cumulative distribution. Note that (14) is the most general definition of conditional cross-covariance that is valid for any statistics, while the Schur complement expression is only justified for Gaussian vectors [9, Ch. 4]. With the goal of deriving an estimator for (14), and following a similar rationale to $\mathring{\bm{\mathsf{u}}}$ and $\mathring{\bm{\mathsf{v}}}$, we define the virtual random variable $\mathring{\mathsf{z}}\triangleq\mathsf{z}_{1}-\mathsf{z}_{2}$, where $\mathsf{z}_{1}$ and $\mathsf{z}_{2}$ are mutually independent and distributed as $\mathsf{z}$. Given that integrating all values of $\mathsf{z}$ is equivalent to doing so for $\mathring{\mathsf{z}}=0$, i.e. $\mathsf{z}_{1}=\mathsf{z}_{2}$, the expectation in (14) can be alternatively expressed as
$\mathbf{C}_{\bm{\mathsf{u}},\bm{\mathsf{v}}|\mathsf{z}}$ $=\int_{\mathbb{R}^{2}}\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})$
$=\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}\int_{\mathbb{R}^{2}}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})=\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0},$ (15)
where $\int_{\mathbb{R}^{2}}\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{1})\operatorname{\mathrm{d}\!}F_{\mathsf{z}}(z_{2})=1$, and since $\mathbf{C}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathring{\mathsf{z}}=0}$ does not depend on the specific values of $z_{1}$ and $z_{2}$ but rather on them being equal. In consequence, conditioning the covariance matrix with respect to $\mathsf{z}$ is equal to conditioning it with respect to $\mathring{\mathsf{z}}=0$. Therefore, the result in (IV) suggests letting the potential confounder data control the pruning of the incomplete U-Statistics in (13). But, since $\mathring{\mathsf{z}}=0$ is an event of zero probability for continuous random variables, the data control should be based on merely small values of $|\mathring{z}|$. Note that the procedure followed here is the same followed in [12].
With the intention of choosing the data pairs for the pruning based on $|\mathring{z}|$, we define the samples of $\mathring{z}$ as follows:
$\mathring{z}_{l,l^{\prime}}\triangleq z(l)-z(l^{\prime}),$ (16)
where $z(l)$ and $z(l^{\prime})$ are i.i.d. samples drawn from $\mathsf{z}$ with $l\neq l^{\prime}$. Then, we let the sorting of $|\mathring{z}_{l,l^{\prime}}|$ (in ascending order) be the one that determines the ordering of the index pairs provided by $\mathrm{f}_{1}(\bm{\cdot})$ and $\mathrm{f}_{2}(\bm{\cdot})$ in (11). Moreover, in view of Remark 1, the amount of pairs $K$ is set to grow as $O(L)$, i.e.
$K_{\alpha}\triangleq\bigl{\lfloor}\tfrac{L\alpha}{2}\bigr{\rfloor},$ (17)
being $1\leq\alpha\ll(L-1)$ a tuning hyper-parameter. While $\alpha=1$ ensures that only very small values of $|\mathring{z}_{l,l^{\prime}}|$ are considered, in the other extreme of $\alpha=L-1$ the U-Statistic becomes complete and there is no conditioning at all, thus encountering the HSIC as a particular case as in (7). A natural trade-off then emerges: while $\alpha$ close to $1$ is desirable for the conditioning idea (IV) to work properly, higher values of $\alpha$ will avoid excessive pruning and provide sufficient statistical accuracy in (13) with $K=K_{\alpha}$. Remarkably, the selection of the hyper-parameter $\alpha$ becomes a minor issue provided that $L$ is sufficiently large, as was shown in [12] under the correlation measure framework between a pair of vectors, and will be seen later on by a numerical example.
IV-A Conditional HSIC
Now that we have determined the sorting and pruning of the data pairs according to the confounder $\mathsf{z}$, let us write a conditional dependence measure as the Frobenius norm of (13):
$\operatorname{\mathrm{tr}}\bigl{(}\widehat{\mathbf{C}}^{\mathrm{H}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\bigr{)}=\tfrac{1}{K_{\alpha}^{2}}\operatorname{\mathrm{tr}}\bigl{(}\mathring{\mathbf{U}}^{\mathrm{H}}\mathring{\mathbf{U}}\mathring{\mathbf{V}}^{\mathrm{H}}\mathring{\mathbf{V}}\bigr{)},$ (18)
where the circularity of the trace has been used. To link the previous expression with kernel-based methods, let us rewrite the zero-mean virtual data matrices $\mathring{\mathbf{U}}$ and $\mathring{\mathbf{V}}$ as follows:
$\mathring{\mathbf{U}}=\tfrac{1}{\sqrt{2}}(\mathbf{U}_{1}-\mathbf{U}_{2}),\qquad\mathring{\mathbf{V}}=\tfrac{1}{\sqrt{2}}(\mathbf{V}_{1}-\mathbf{V}_{2}),$ (19)
where
$\mathbf{U}_{a}\triangleq[\mathbf{u}_{\mathrm{f}_{a}(1)},\dots,\mathbf{u}_{\mathrm{f}_{a}(K_{\alpha})}],\;\;\mathbf{V}_{a}\triangleq[\mathbf{v}_{\mathrm{f}_{a}(1)},\dots,\mathbf{v}_{\mathrm{f}_{a}(K_{\alpha})}],$ (20)
for $a=\{1,2\}$. Accordingly, (18) is then rewritten as
$\operatorname{\mathrm{tr}}\bigl{(}\widehat{\mathbf{C}}^{\mathrm{H}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\widehat{\mathbf{C}}_{\mathring{\bm{\mathsf{u}}},\mathring{\bm{\mathsf{v}}}|\mathsf{z}}\bigr{)}$ (21)
$=\tfrac{1}{4K_{\alpha}^{2}}\operatorname{\mathrm{tr}}\bigl{(}(\mathbf{U}_{1}-\mathbf{U}_{2})^{\mathrm{H}}(\mathbf{U}_{1}-\mathbf{U}_{2})(\mathbf{V}_{1}-\mathbf{V}_{2})^{\mathrm{H}}(\mathbf{V}_{1}-\mathbf{V}_{2})\bigr{)}.$
Taking the limit $M\rightarrow\infty$, kernel matrices are obtained from the products among $\mathbf{U}$ and $\mathbf{V}$ in (21):
$\mathbf{K}_{a,a^{\prime}}\triangleq\lim_{M\rightarrow\infty}\mathbf{U}_{a}^{\mathrm{H}}\mathbf{U}_{a^{\prime}}$ $=\mathbf{S}\mathopen{}\mathclose{{}\left(\mathbf{x}_{a},\mathbf{x}_{a^{\prime}}}\right)\quad$ (22a)
$\mathbf{Q}_{a,a^{\prime}}\triangleq\lim_{M\rightarrow\infty}\mathbf{V}_{a}^{\mathrm{H}}\mathbf{V}_{a^{\prime}}$ $=\mathbf{S}\mathopen{}\mathclose{{}\left(\mathbf{y}_{a},\mathbf{y}_{a^{\prime}}}\right),$ (22b)
where $\mathbf{S}\mathopen{}\mathclose{{}\left(\bm{\cdot},\bm{\cdot}}\right)$ is given in (9). Therefore, the new relevant data subsets for inferring conditional dependence are given by
$[\mathbf{x}_{a}]_{k}\triangleq x(f_{a}(k)),\quad[\mathbf{y}_{a}]_{k}\triangleq y(f_{a}(k)).$ (23)
Finally, the resulting C-HSIC can be expressed as follows:
[…]
with $\breve{\mathbf{K}}\triangleq\mathbf{K}_{1,1}+\mathbf{K}_{2,2}-\mathbf{K}_{1,2}-\mathbf{K}_{2,1}$ and $\breve{\mathbf{Q}}\triangleq\mathbf{Q}_{1,1}+\mathbf{Q}_{2,2}-\mathbf{Q}_{1,2}-\mathbf{Q}_{2,1}$. Note that, as a result of the U-Statistics implementation, each entry of the new kernel-based matrices involves four data samples of the same source ($4$-tuples), in contrast to only the pairs typically involved in classical kernel methods. This fact, along with the lack of $\mathbf{P}$, are the main distinctive features of the C-HSIC (24) vs. HSIC (7).
V Numerical illustration
In order to test the proposed method, we aim at generating uncorrelated data with a controlled amount of co-information [17] (also called interaction information [18]), defined as
$\mathrm{I}(\mathsf{x};\mathsf{y};\mathsf{z})=\mathrm{I}(\mathsf{x};\mathsf{y})-\mathrm{I}(\mathsf{x};\mathsf{y}|\mathsf{z}),$ (25)
such that the potentials of the method can be clearly highlighted. Similar ideas on modeling conditional dependences can be found in [19]. Since co-information can be either positive or negative, two scenarios are studied: $\mathcal{M}^{+}$ for $\mathrm{I}(\mathsf{x};\mathsf{y};\mathsf{z})>0$ and $\mathcal{M}^{-}$ for $\mathrm{I}(\mathsf{x};\mathsf{y};\mathsf{z})<0$, whose variables are as follows:
+:$\mathcal{M}^{+}:\begin{cases}\mathsf{x}=\sqrt{\gamma}\mathsf{a}\mathsf{p}+\mathsf{v}\\ \mathsf{y}=\sqrt{\gamma}\mathsf{a}\mathsf{q}+\mathsf{w}\\ \mathsf{z}=\mathsf{a}\end{cases},\hskip 5.0pt\mathcal{M}^{-}:\begin{cases}\mathsf{x}=\sqrt{\gamma}\mathsf{b}\mathsf{p}+\mathsf{v}\\ \mathsf{y}=\sqrt{\gamma}\mathsf{c}\mathsf{q}+\mathsf{w}\\ \mathsf{z}=\mathsf{b}-\mathsf{c}\end{cases}.$ (26)
The internal i.i.d. random variables are distributed as $\mathsf{a},\mathsf{b},\mathsf{c}\sim\mathcal{U}(0,\sqrt{3\text{}})$ (uniform), $\mathsf{v},\mathsf{w}\sim\mathcal{N}(0,1)$ (normal), and $\mathsf{p},\mathsf{q}\sim\mathrm{Bern}_{1/2}\{-1,1\}$ (equiprobable Bernoulli). Parameter $\gamma$ is the signal-to-noise ratio associated to the measurements and controls the total amount of absolute co-information. In model $\mathcal{M}^{+}$, $\mathsf{x}$ and $\mathsf{y}$ are dependent due to the influence of $\mathsf{a}$ in both, but they are conditionally independent, since knowing $\mathsf{z}$ (i.e. $\mathsf{a}$) implies that $\mathsf{x}$ and $\mathsf{y}$ become solely driven by independent phenomena ($\mathsf{v}$ and $\mathsf{w}$). By contrast, $\mathsf{x}$ and $\mathsf{y}$ are marginally independent in model $\mathcal{M}^{-}$, but they become conditionally dependent, since the knowledge of $\mathsf{z}$ correlates the possible joint values of $\mathsf{b}$ and $\mathsf{c}$. In both models, $\mathsf{x}$, $\mathsf{y}$ and $\mathsf{z}$ are mutually uncorrelated due to the multiplicative effect of mutually independent variables $\mathsf{p}$ and $\mathsf{q}$, so correlation measures are unable to discover any data association.
Regarding the choice of kernel, the universal Gaussian kernel is used, which yields $\kappa(s)=\exp(-(\frac{s}{\hat{\sigma}L^{-1/5}})^{2})$. Note that the kernel length scales in proportion with the sample standard deviation $\hat{\sigma}$ of the data and in inverse proportion with the data size $L$. The reader is referred to [7, Appendix D][20, 21] for a justification of the typical power law $O(L^{-1/5})$.
Figure 1: Model $\mathcal{M}^{+}$. Orange: $L=100$; blue: $L=600$. Markers denote the empirical average while bands indicate the standard deviation. (a)-cross marker: HSIC; (a)-circle marker: C-HSIC; (b): C-HSIC with random pruning.
Figure 2: Model $\mathcal{M}^{-}$. Orange: $L=100$; blue: $L=600$. Markers denote the empirical average while bands indicate the standard deviation. (a)-cross marker: HSIC; (a)-circle marker: C-HSIC; (b): C-HSIC with random pruning.
In Fig. 1 and Fig. 2, the measures of dependence are shown as a function of $\gamma$ for both models with a moderate hyper-parameter $\alpha=4$ from (17). This yields $8\%$ of the total data pairs for $L=100$ and $1.3\%$ for $L=600$ (and even less for higher $L$). For model $\mathcal{M}^{+}$ (Fig. 1), the C-HSIC correctly depicts small conditional dependence, while the HSIC confirms that marginal dependence is high for moderate $\gamma$ values. Conversely, Fig. 2 exhibits the capability of the C-HSIC to discover significant conditional dependence for moderate $\gamma$ values for model $\mathcal{M}^{-}$, while marginal dependence is confirmed to be small by the HSIC at any value of $\gamma$.
To get further insights on the proposed ideas, Fig. 1-(b) and Fig. 2-(b) show the performance of the C-HSIC when the pruning of data pairs in the U-Statistics is random, which means that the pair $(l(k),l^{\prime}(k))$ is not controlled by the confounder. It is seen that the C-HSIC produces a measure that mimics the marginal dependence one, thus confirming the main property of unbiasedness related to the general U-Statistics theory, and also showing its consequent increase in variance. It is also important to note that the selected value of $\alpha$, and the small percentage of data that yields this choice, confirms that the increase in variance induced by the pruning becomes small for sufficiently large data sizes due to Remark 1, making $\alpha$ a noncritical free hyper-parameter.
VI Conclusions
The classical HSIC measure for marginal statistical dependence has been reinterpreted as a correlation measure on a finite but high-dimensional space based on windowed steering vectors. From the original mapping, we can establish an insightful connection with kernels by letting the dimension grow to infinity. Instead of being merely a reinterpretation of the HSIC, this vision demonstrates its main power under the more involved conditional dependence framework. In particular, the illustrated identification of the HSIC as a sample covariance estimator opens the possibility of leveraging U-Statistics for this task. Thanks to this formulation, we can provide a novel route for performing statistical conditioning by pruning data pairs based on the pairwise differences of the confounder under suspicion. Furthermore, the proposed measure of conditional dependence does not require matrix inversions, which has the advantage of reduced computational complexity and the avoidance of addressing ill-conditioned matrices. Further work should study the potential of the proposed method with richer data sets.